# Predicting Drug-Gene Relations via Analogy Tasks with Word Embeddings

### v2

@@ Acknowledgements
This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3, JST SPRING JPMJSP2110.
@@
H.Y., R.H., K.A., Ma.O. and H.S. conceived the experiments, H.Y., R.H., K.A., and K.M. conducted the experiments, H.Y., R.H., K.A., K.M., S.S., Ma.O., and H.S. analysed the results, R.H. and Mo.O. surveyed existing research. All authors reviewed the manuscript.
@@ 1.2.1 Global setting by year
Similar to $\hat{\mathbf{v}}$ in Eq. (7), to measure the performance of the estimator $\hat{\mathbf{v}}^{y}$ in Eq. (20), we prepare the evaluation of the analogy tasks. Using the projection operations in Eq. (10), we define $D^{y}:=\pi_{\mathcal{D}^{y}}(\mathcal{R}^{y})\subset\mathcal{D}^{y}$ and $G^{y}:=\pi_{\mathcal{G}^{y}}(\mathcal{R}^{y})\subset\mathcal{G}^{y}$.
@@ 1.2.2 Global setting to predict unknown relations by year
where $\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}^{y\mid L_{y}}$. We define the estimator $\hat{\mathbf{v}}^{y\mid L_{y}}$ in Eq. (S4) by using $\mathcal{R}^{y\mid L_{y}}$ instead of $\mathcal{R}^{y}$ in the estimator $\hat{\mathbf{v}}^{y}$ in Eq. (20). For easy comparison with Eq. (5), Eq. (S4) is rewritten as the difference of mean vectors:
@@
$\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{d}.$ (S6)
@@
$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{d\in\mathcal{D}_{p}\cap D}\{|[d]|\}\}$ 16.346 16.302
$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in\mathcal{G}_{p}\cap G}\{|[g]|\}\}$ 101.765 91.488
@@ 1.2.3 Pathway-wise setting by year
$\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y}}\mathbf{u}_{d}.$ (S12)
@@ 1.2.4 Pathway-wise setting to predict unknown relations by year
$\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}}\mathbf{u}_{d}.$ (S18)
@@ G $d\in D$ $[d]$ $\mathcal{G}$
1.4 Embeddings
The hyperparameters used to train our skip-gram are shown in Table S4.
1.5 Datasets
1.5.1 Conversion from MeSH ID to drug name
@@
BioConceptVec ${}^{{}^{\prime}}$ Random 0.000 0.001 0.001
@@
Our embeddings ${}^{{}^{\prime}}$ Random 0.000 0.003 0.002
@@ 2.1.3 Predictions by GPT models
The prediction results from the GPT models are shown in Table S9. The set of query drugs and the set of correct target genes depend not only on the settings G, P1, and P2 but also on the vocabulary set. Here, we conducted experiments using vocabularies obtained from BioConceptVec or our skip-gram embeddings. The best accuracies for all combinations of vocabularies and settings are summarized in the Table 3 in the main text.

### v4

@@ Acknowledgements
This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3.
@@
H.Y., R.H., K.A., Ma.O. and H.S. conceived the experiments, H.Y., R.H., K.A., Y.Z. and K.M. conducted the experiments, H.Y., R.H., K.A., K.M., S.S., Y.Z., Ma.O., and H.S. analysed the results, R.H. and Mo.O. surveyed existing research. All authors reviewed the manuscript.
@@ 1.1.1 Pathway-wise setting
For easy comparison with Eq. (5), Eq. (14) is rewritten as the difference of mean vectors:
$\hat{\mathbf{v}}_{p}=\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{d}\},$ (S1)
$\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{d}\}=\frac{1}{|\mathcal{R}_{p}|}\sum_{(d,g)\in\mathcal{R}_{p}}\mathbf{u}_{d},\quad\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{g}\}=\frac{1}{|\mathcal{R}_{p}|}\sum_{(d,g)\in\mathcal{R}_{p}}\mathbf{u}_{g}.$ (S2)
@@ 1.2.1 Global setting by year
Similar to the estimator $\hat{\mathbf{v}}$ in Eq. (7), we define an estimator $\hat{\mathbf{v}}^{y}$ for the relation vector $\mathbf{v}^{y}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}^{y}$:
$\hat{\mathbf{v}}^{y}:=\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}^{y}\right|}\sum_{(d,g)\in\mathcal{R}^{y}}(\mathbf{u}_{g}-\mathbf{u}_{d}),$ (S5)
where $\mathrm{E}_{\mathcal{R}^{y}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}^{y}$. For easy comparison with Eq. (5), Eq. (S5) is rewritten as the difference of mean vectors:
$\hat{\mathbf{v}}^{y}=\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{d}\},$ (S6)
$\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}^{y}|}\sum_{(d,g)\in\mathcal{R}^{y}}\mathbf{u}_{d},\quad\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}^{y}|}\sum_{(d,g)\in\mathcal{R}^{y}}\mathbf{u}_{g}.$ (S7)
Similar to $\hat{\mathbf{v}}$ in Eq. (7), to measure the performance of the estimator $\hat{\mathbf{v}}^{y}$ in Eq. (S5), we prepare the evaluation of the analogy tasks. Using the projection operations in Eq. (10), we define $D^{y}:=\pi_{\mathcal{D}^{y}}(\mathcal{R}^{y})\subset\mathcal{D}^{y}$ and $G^{y}:=\pi_{\mathcal{G}^{y}}(\mathcal{R}^{y})\subset\mathcal{G}^{y}$.
@@ 1.2.2 Global setting to predict unknown relations by year
where $\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}^{y\mid L_{y}}$. We define the estimator $\hat{\mathbf{v}}^{y\mid L_{y}}$ in Eq. (S9) by using $\mathcal{R}^{y\mid L_{y}}$ instead of $\mathcal{R}^{y}$ in the estimator $\hat{\mathbf{v}}^{y}$ in Eq. (S5). For easy comparison with Eq. (5), Eq. (S9) is rewritten as the difference of mean vectors:
@@
$\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{d},\quad\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{g}.$ (S11)
@@
$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{d\in\mathcal{D}_{p}\cap D}\{|[d]|\}\}$ 2.738 2.776
$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in\mathcal{G}_{p}\cap G}\{|[g]|\}\}$ 7.131 6.714
@@ 1.2.3 Pathway-wise setting by year
$\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y}}\mathbf{u}_{d},\quad\mathrm{E}_{\mathcal{R}_{p}^{y}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y}}\mathbf{u}_{g}.$ (S17)
@@ 1.2.4 Pathway-wise setting to predict unknown relations by year
$\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}}\mathbf{u}_{d},\quad\mathrm{E}_{\mathcal{R}_{p}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}_{p}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}_{p}^{y\mid L_{y}}}\mathbf{u}_{g}.$ (S23)
@@ G $d\in D$ $[d]$ $\mathcal{G}$
1.4 Embeddings
Figure S2 shows the distribution of the sizes of the answer sets for each drug $d$ in settings G, P1, and P2 for BioConceptVec and our skip-gram embeddings. Similarly, Fig. S3 shows the distribution of the sizes of the answer sets for each gene $g$ in settings ${}^{{}^{\prime}}$, ${}^{{}^{\prime}}$, and ${}^{{}^{\prime}}$ for BioConceptVec and our skip-gram embeddings. The hyperparameters used to train our skip-gram are shown in Table S4.
1.5 Datasets
1.5.1 Conversion from MeSH ID to drug name
(a) BioConceptVec
(b) Our embeddings
Figure S2: Distribution of the sizes of the answer sets for each drug $d$ in the settings for predicting genes from drugs. The mean values are also shown: for setting G, the value of $\mathrm{E}_{d\in D}\{|[d]|\}$ from Table 2; for setting P1, the value of $\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{d\in D_{p}}\{|[d]_{p}|\}\}$ from Table S1; and for setting P2, the value of $\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{d\in\mathcal{D}_{p}\cap D}\{|[d]|\}\}$ from Table S1.
(a) BioConceptVec
(b) Our embeddings
Figure S3: Distribution of the sizes of the answer sets for each gene $g$ in the settings for predicting drugs from genes. The mean values are also shown: for setting ${}^{{}^{\prime}}$, the value of $\mathrm{E}_{g\in G}\{|[g]|\}$ from Table 2; for setting ${}^{{}^{\prime}}$, the value of $\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in G_{p}}\{|[g]_{p}|\}\}$ from Table S1; and for setting ${}^{{}^{\prime}}$, the value of $\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in\mathcal{G}_{p}\cap G}\{|[g]|\}\}$ from Table S1.
@@ 1.5.1 Conversion from MeSH ID to drug name
BioConceptVec ${}^{{}^{\prime}}$ Random 0.011 0.075 0.036
@@
Our embeddings ${}^{{}^{\prime}}$ Random 0.010 0.078 0.037
@@ 1.5.3 Conversion from KEGG ID to MeSH ID
1.6 Details of predictions using random baseline
This section describes the sets used for sampling in the random baseline across different settings. For settings G, P1, and P2, the set used was $G$. Similarly, $D$ was used for ${}^{{}^{\prime}}$, ${}^{{}^{\prime}}$, and ${}^{{}^{\prime}}$, while $G^{y}$ was used for settings Y1, P1Y1, and P2Y1. For settings Y2, P1Y2, and P2Y2, $G^{y\mid L_{y}}\cup G^{y\mid U_{y}}$ was used instead of $G^{y\mid U_{y}}$; This was intended to prevent extremely high random baseline performance when the size of $G^{y\mid U_{y}}$ is very small.
@@ 1.7 Details of predictions using the GPT series
1.8 Details of predictions by TransE
This section describes the details of the experimental settings for TransE. Embeddings are learned for each of settings G, P1, and P2. In our experiments, we utilized TransE with the following hyperparameters: a batch size of 1024, embedding dimensions set to 500, a learning rate of 0.0001, and a total of 800 iteration steps. Additionally, we adopted the self-adversarial negative sampling method from RotatE [60], configuring the negative sampling size to 512 and the sampling temperature to 0.5. For each setting, the triplets of drug $d$, gene $g$, and drug-gene relation $(d,g)$ are used to evaluate performance in analogy tasks and are then split into 60% training, 20% validation, and 20% test data. In setting G, only a single type of relation, the drug-gene relation, is considered. In settings P1 and P2, the drug-gene relation $(d,g)$ is distinguished based on each pathway $p\in\mathcal{P}$, resulting in $|\mathcal{P}|$ different relations. For entities in the validation and test data that do not appear in the training data, random embeddings are simply used. Since KGE typically uses the entire set of head and tail entities as its search space, we similarly define the search space for TransE as the entire set of drugs and genes that have drug-gene relations for each setting. Additionally, note that in Table 3, while random baseline, GPT series, and our method evaluate performance on the entire dataset, TransE evaluates performance on the 20% test data.
@@ 2.1.3 Predictions by GPT models
The prediction results from the GPT models are shown in Table S9. The set of query drugs and the set of correct target genes depend not only on settings G, P1, and P2 but also on the vocabulary set. Here, we conducted experiments using vocabularies obtained from BioConceptVec or our skip-gram embeddings. The best accuracies for all combinations of vocabularies and settings are summarized in the Table 3 in the main text.
@@ 2.2.2 Pahtway-wise setting by year
2.3 Correlation between answer set sizes and predicted ranks
(a) BioConceptVec
(b) Our embeddings
Figure S5: Scatter plots of the answer set sizes for each drug $d$ and the ranks of the predicted genes.
Fig. S5 shows scatter plots and correlation coefficients between the answer set sizes for each drug $d$ and the ranks of the predicted genes obtained by adding relation vectors in settings G, P1, and P2. Since smaller answer set sizes are more frequent (see Fig. S2 in Supplementary Information 1.4 for the distribution of answer set sizes), we calculated a weighted Pearson correlation coefficient, where the weight is the reciprocal of the frequency of each answer set size.
2.4 Details of comparison of analogy computation and TransE
This section provides details regarding the experiments presented in Fig. 4, comparing our analogy computation with TransE. We conducted experiments under setting P1, which is a simple setting and involves multiple drug-gene relations. For drug-gene relations in Setting P1, we fixed 20% of the data as validation data and another 20% as test data, following the same setting for TransE in Table 3, and used them consistently across both methods. From the remaining 60% of the data, we created training subsets ranging from 10% to 60%, ensuring that each smaller subset is contained within the next larger subset. Formally, if we define the 10% to 60% training sets as $\mathcal{T}_{10},\ldots,\mathcal{T}_{60}$, then $\mathcal{T}_{10}\subset\mathcal{T}_{20}\subset\mathcal{T}_{30}\subset\mathcal{T}_{40}\subset\mathcal{T}_{50}\subset\mathcal{T}_{60}$.
We explain the experimental settings of our method. We used our skip-gram embeddings. At test time, for pathways present in the training data, we used the relation vector $\hat{\mathbf{v}}_{p}$ defined in Eq. (14). For pathways not present in the training data, we used the relation vector $\hat{\mathbf{v}}$ computed from all drug-gene relations in the training data, as defined in Eq. (7). The search space consisted of all genes appearing in the training, validation, and test data.
The experimental settings for TransE follow those described in Supplementary Information 1.8.